# Symbiotic Game and Foundation Models for Cyber Deception Operations in Strategic Cyber Warfare

### v1

@@ 3.4 Foundation models for Mechanism Design for Security Games
Many components of interactions are amenable to design or control, including incentives (player payoffs) [43], game dynamics (course direction shaped by actions), and information disclosed to players during gameplay [92, 98]. The design process itself can take either an offline or online approach. Offline design involves training FMs to ascertain optimal game configurations, while online design employs a reinforcement learning process that dynamically adapts to player actions. Both approaches leverage hyper-combinations of analytics through FMs or the design of hyper-FMs to achieve their objectives.
@@ 4 Cyber Deception Game and Foundation Models
There are several quintessential domains in which FMs can play a significant role in cyber deception. Cyber deception operations involve numerous tactical components, including information gathering, detection and response, configuration and deployment of honeypots, and engagement with attacks. In this section, we specifically discuss the connection between FMs and GMs in the use of defensive cyber deception for attack engagement and defensive response. Once an attacker is in the network, the goal of defensive cyber deception is to thwart and mislead the attacker who are attempting to breach or infiltrate the network or systems by deploying techniques such as fake credentials and honeypots to deceive and divert attackers’ attention away from valuable assets and towards simulated or less critical targets. Many game-theoretic approaches have been proposed to describe various scenarios, [77, 53]. For instance, a signaling game framework has been used to capture the asymmetric information between the two players [76]. In [39], a one-sided information stochastic game is used to examine the effects of deception on the attacker’s belief. The study has explored the sequential nature of attacks and investigates how an attacker’s beliefs evolve, influencing their actions. It has demonstrated strategies for defenders to manipulate attacker beliefs effectively, hindering attackers from achieving their objectives and minimizing network damage.
@@ 4.1 Challenges in Cyber Deception
One key challenge inherent in cyber deception lies in understanding attacker behaviors through proactive engagements. It is crucial to deduce the incentives and motivations driving attackers by leveraging heterogeneous sources of data, including traffic data, log data, and event data. By doing so, defenders can gain insight into TTPs used by the attackers and their objectives. It enables the defender to prescribe the most effective responses to emerging threats.
@@
The following is an example of the parsing of PCAP data that captures the scenario when an attacker responds to the patching of network systems. Their behavior may change as they adapt to the evolving security measures. The attacker initially attempts to exploit known vulnerabilities in the target system. However, as the target system patches these vulnerabilities, the attacker’s exploit attempts become unsuccessful. The attacker responds by continuing to probe the target for other potential vulnerabilities and, upon identifying a successful attack vector, exfiltrates data to an external server. This PCAP table illustrates an attacker’s adaptive behavior in response to network system patching, highlighting their persistence and ongoing efforts to exploit vulnerabilities and compromise the target environment.
[…]
90 Frame Action Source Destination Protocol Flags 1 SYN packet sent Attacker_MAC Target_MAC TCP SYN 2 SYN-ACK response Target_MAC Attacker_MAC TCP SYN 3 Exploitation attempt Attacker_MAC Target_MAC TCP PSH ACK 4 Exploit attempt unsuccessful Target_MAC Attacker_MAC TCP RST 5 SYN packet sent Attacker_MAC Target_MAC TCP SYN 6 SYN-ACK response Target_MAC Attacker_MAC TCP SYN ACK 7 Data exfiltration attempt Attacker_MAC External_Server_MAC TCP PSH ACK 8 Data exfiltration successful External_Server_MAC Attacker_MAC TCP ACK
@@
CAPTION: The presented PCAP table shows the adaptive nature of an attacker encountering patched network systems. Initially, the attacker seeks to exploit known vulnerabilities within the target system. However, as the target’s vulnerabilities are patched, the attacker’s exploit attempts are thwarted. Undeterred, the attacker adapts by persistently probing the target for alternative vulnerabilities. Upon discovering a successful attack vector, the attacker shifts tactics, opting to exfiltrate data to an external server.
@@
FMs play a pivotal role in facilitating the learning process and empowering defenders to navigate the complexities of cyber threats effectively. By harnessing FM’s capabilities, defenders can gain valuable insights into the attacker’s knowledge and intentions, enabling them to map out the cyber landscape and identify potential vulnerabilities. Various types of FMs serve distinct purposes in the intrusion response process. For instance, LLMs excel at parsing and comprehending the nuanced nuances of the attacker’s intent, enabling defenders to decipher malicious communications and identify potential threats. Similarly, Decision Transformers [52, 20, 57] leverage sophisticated algorithms and reinforcement learning techniques to predict attacker behaviors and make optimal decisions in real time. By analyzing patterns and trends in attacker activity, Decision Transformers empower defenders to anticipate threats and implement proactive defense measures.
@@
Apart from supporting the ability of the defender to use his knowledge to predict the attacker’s paths, FMs can also be used to allow the defender to synthesize and acquire knowledge over time. Understanding the attacker’s prior behavior and potential scenarios is crucial for the defender to quickly update its security policies. Proficient knowledge enables the defender to adapt promptly and avoid extensive trial-and-error efforts. This knowledge is typically encoded using symbolic representations of rules, laws, or constraints. For instance, rules may be expressed through first-order or temporal logic statements, while laws can be encapsulated in rule-of-thumb equations, offering approximate responses. Constraints capture feasible directions for updates, which can be weighted to prioritize specific directions. Acquiring knowledge presents a challenge. One immediate method involves knowledge sharing, where insights on unknown attacks can be gleaned from those who have encountered similar incidents. In [27, 102, 103], mechanisms for knowledge sharing among intrusion detection systems are explored, introducing an incentive-compatible approach to encourage data sharing and expedite responses to zero-day attacks. Alternatively, knowledge can be synthesized from individual experiences and data. FMs play a pivotal role here, encoding experiences and data into symbolic representations conducive to formal reasoning, computation, and optimization of security policies. Recent works in [89, 90] provide promising approaches of using FMs to synthesize knowledge from datasets.
@@
Another challenging area of cyber deception is the configuration and utilization of honeypots in engaging attackers to acquire information and knowledge. Honeypot games have been used to formalize the deployment and configuration decisions by taking into account the adversarial interactions and the engagement goals. For instance, in [44], a reinforcement learning algorithm is proposed to map out an attacker’s behavior. The engagement has to interact with the attacker, making sure that the attacker does not exist in the honeynetwork. By observing and analyzing attacker interactions with honeypots, defenders gain valuable insights into emerging threats and attacker behaviors. There is a tradeoff between learning the attacker’s behaviors and removing the attacker directly. The knowledge of attackers can be useful to further protect ourselves from zero-day attacks in the future. However, removing the attackers directly can lead to fewer risks.
@@ 4.2 A Neurosymbolic Learning Approach to Cyber Deception
FMs serve as pivotal tools across various functions including reinforcement learning, knowledge assimilation, formation of conjectures, and contextual representation. Moreover, they facilitate the coordination of multiple components in the neurosymbolic learning. Figure 7 further illustrates the synergy between FMs and GMs for cyber deception. A sequence of tactical games $G_{1},G_{2},G_{3}$ unfold between an attacker and a defender. The defender leverages $F_{1}$ to conjecture the attacker’s behavior and utilizes $F_{2}$ to adapt security measures based on acquired observations. This conjecture and subsequent updates stem from the defender’s knowledge, which is symbolically represented using $F_{4}$.
@@
Figure 7: Interconnections between FMs and GMs for cyber deception: The tactical games are played between an attacker and a defender. A defender uses $F_{1}$ to conjecture the attacker’s behavior and uses $F_{2}$ to update the security knowledge. The conjecture and the update are based on the knowledge of the defender. The knowledge is symbolically represented and captured using $F_{4}$. The conjectured learning is driven by $F_{5}$, which coordinates the conjecture update and the policy update. The tactical game is sequentially played over time to achieve its operational-level objective. The changes in the environment and the operational-level objective can lead to the adaptation of tactical-level games. This change is represented by the contextual parameter $\theta$, which requires $F_{5}$ to process heterogeneous data sources to identify the context and $F_{6}$ to coordinate between the tactical level games and the contextual changes. Each context has an associated knowledge set $K$, which updates itself when the context changes. FM $F_{7}$ provides a set of rules for policy updates and conjecture formation under a context. FM $F_{4}$ is used to figure out the set of relevant knowledge which $F_{7}$ can choose from. $F_{4}$ can also create new rules for new contexts transferrable from other contexts or learned from datasets.
@@ 4.3 Emerging Challenges
Another practical challenge arises from the inference time current FMs consume for outputting the next token. In the presence of advanced adversarial attacks, and in particular, cyber threats, real-time inference is the prerequisite of a rapid defense response to counteract attackers’ swift moves, for example, lateral movement within the system. Accelerating the FM’s inference with long contexts is imperative for predictive and prescriptive security analytics.

### v2

@@ 3.4 Foundation models for Mechanism Design for Security Games
Many components of interactions are amenable to design or control, including incentives (player payoffs) Huang et al. (2022), game dynamics (course direction shaped by actions), and information disclosed to players during gameplay Yang et al. (2023); Zhang and Zhu (2023). The design process itself can take either an offline or online approach. Offline design involves training FMs to ascertain optimal game configurations, while online design employs a reinforcement learning process that dynamically adapts to player actions. Both approaches leverage hyper-combinations of analytics through FMs or the design of hyper-FMs to achieve their objectives.
@@ 4 Cyber Deception Game and Foundation Models
There are several quintessential domains in which FMs can play a significant role in cyber deception. Cyber deception operations involve numerous tactical components, including information gathering, detection and response, configuration and deployment of honeypots, and engagement with attacks. In this section, we specifically discuss the connection between FMs and GMs in the use of defensive cyber deception for attack engagement and defensive response. Once an attacker is in the network, the goal of defensive cyber deception is to thwart and mislead the attacker who are attempting to breach or infiltrate the network or systems by deploying techniques such as fake credentials and honeypots to deceive and divert attackers’ attention away from valuable assets and towards simulated or less critical targets. Many game-theoretic approaches have been proposed to describe various scenarios, Pawlick et al. (2019); Kamhoua et al. (2021). For instance, a signaling game framework has been used to capture the asymmetric information between the two players Pawlick et al. (2018). In Horák et al. (2017), a one-sided information stochastic game is used to examine the effects of deception on the attacker’s belief. The study has explored the sequential nature of attacks and investigates how an attacker’s beliefs evolve, influencing their actions. It has demonstrated strategies for defenders to manipulate attacker beliefs effectively, hindering attackers from achieving their objectives and minimizing network damage.
@@ 4.1 Challenges in Cyber Deception
One key challenge inherent in cyber deception lies in understanding attacker behaviors through proactive engagements. It is crucial to infer the incentives and motivations driving attackers by leveraging heterogeneous sources of data, including traffic data, log data, and event data. By doing so, defenders can gain insight into TTPs used by the attackers and their objectives. It enables the defender to prescribe the most effective responses to emerging threats.
@@
The following Table 1 is an example of the parsing of PCAP data that captures the scenario when an attacker responds to the patching of network systems and aims for data exfiltration. Their behavior may change as they adapt to the evolving security measures. The attacker initially attempts to exploit known vulnerabilities in the target system. However, as the target system patches these vulnerabilities, the attacker’s exploit attempts become unsuccessful. The attacker responds by continuing to probe the target for other potential vulnerabilities and, upon identifying a successful attack vector, exfiltrates data to an external server. This PCAP table illustrates an attacker’s adaptive behavior in response to network system patching, highlighting their persistence and ongoing efforts to exploit vulnerabilities and compromise the target environment.
Frame Action Source Destination Protocol Flags
1 SYN packet sent Attacker_MAC Target_MAC TCP SYN
2 SYN-ACK response Target_MAC Attacker_MAC TCP SYN
3 Exploitation attempt Attacker_MAC Target_MAC TCP PSH ACK
4 Exploit attempt unsuccessful Target_MAC Attacker_MAC TCP RST
5 SYN packet sent Attacker_MAC Target_MAC TCP SYN
6 SYN-ACK response Target_MAC Attacker_MAC TCP SYN ACK
7 Data exfiltration attempt Attacker_MAC External_Server_MAC TCP PSH ACK
8 Data exfiltration successful External_Server_MAC Attacker_MAC TCP ACK
@@
Table 1: The presented PCAP table shows the adaptive nature of an attacker encountering patched network systems. Initially, the attacker seeks to exploit known vulnerabilities within the target system. However, as the target’s vulnerabilities are patched, the attacker’s exploit attempts are thwarted. Undeterred, the attacker adapts by persistently probing the target for alternative vulnerabilities. Upon discovering a successful attack vector, the attacker shifts tactics, opting to exfiltrate data to an external server.
Frame Action Source Destination Protocol Flags
1 Reconnaissance by hesitant attacker Attacker_IP Network_Router_IP TCP SYN
2 SYN packet forwarded to honeypot response Attacker_IP Honeypot_IP TCP SYN
3 Honeypot responds with SYN-ACK Honeypot_IP Attacker_IP TCP SYN ACK
4 Attacker sends ACK to establish connection Attacker_IP Honeypot_IP TCP ACK
5 Honeypot acknowledges connection Honeypot_IP Attacker_IP TCP ACK
6 Attacker probes honeypot cautiously Attacker_IP Honeypot_IP TCP PSH ACK
7 Honeypot responds to attacker probing Honeypot_IP Attacker_IP TCP PSH ACK
8 Attacker engages further with honeypot Attacker_IP Honeypot_IP TCP PSH ACK
Table 2: The table illustrates a strategic interaction between a cautious attacker and a honeypot deployed by the defender. The attacker initially conducts a cautious scan of the network for potential targets, exhibiting hesitant behavior. The honeypot strategically responds to the attacker’s SYN packet with a SYN-ACK, mimicking a legitimate service to attract the attacker. The attacker acknowledges the honeypot’s SYN-ACK by sending an ACK packet to establish a connection. Subsequently, the attacker proceeds cautiously, sending probing packets to the honeypot to gather information without revealing their true intentions. The honeypot responds to the attacker’s probing, maintaining the illusion of a legitimate service and engaging in interactive communication.
Another example of the attacker’s behavior is illustrated using the following PCAP data in Table 2. The attacker interacts with honeypots in a discretionary manner. The attacker initially exhibits hesitant behavior, cautiously interacting with the honeypot to gather information and assess the situation. The honeypot, strategically placed by the defender, maintains the illusion of a legitimate service, effectively attracting and engaging the attacker without raising suspicion. This strategic approach allows the defender to gather valuable intelligence about the attacker’s behavior and intentions.
@@ 8 Attacker engages further with honeypot Attacker_IP Honeypot_IP TCP PSH ACK
FMs play a pivotal role in facilitating the learning process and empowering defenders to navigate the complexities of cyber threats effectively. By harnessing FM’s capabilities, defenders can gain valuable insights into the attacker’s knowledge and intentions, enabling them to map out the cyber landscape and identify potential vulnerabilities. Various types of FMs serve distinct purposes in the intrusion response process. For instance, LLMs excel at parsing and comprehending the nuanced nuances of the attacker’s intent, enabling defenders to decipher malicious communications and identify potential threats. Similarly, Decision Transformers Janner et al. (2021); Chen et al. (2021); Li et al. (2023) leverage sophisticated algorithms and reinforcement learning techniques to predict attacker behaviors and make optimal decisions in real time. By analyzing patterns and trends in attacker activity, Decision Transformers empower defenders to anticipate threats and implement proactive defense measures.
@@
Apart from supporting the ability of the defender to use his knowledge to predict the attacker’s paths, FMs can also be used to allow the defender to synthesize and acquire knowledge over time. Understanding the attacker’s prior behavior and potential scenarios is crucial for the defender to quickly update its security policies. Proficient knowledge enables the defender to adapt promptly and avoid extensive trial-and-error efforts. This knowledge is typically encoded using symbolic representations of rules, laws, or constraints. For instance, rules may be expressed through first-order or temporal logic statements, while laws can be encapsulated in rule-of-thumb equations, offering approximate responses. Constraints capture feasible directions for updates, which can be weighted to prioritize specific directions. Acquiring knowledge presents a challenge. One immediate method involves knowledge sharing, where insights on unknown attacks can be gleaned from those who have encountered similar incidents. In Fung et al. (2011); Zhu et al. (2011, 2012), mechanisms for knowledge sharing among intrusion detection systems are explored, introducing an incentive-compatible approach to encourage data sharing and expedite responses to zero-day attacks. Alternatively, knowledge can be synthesized from individual experiences and data. FMs play a pivotal role here, encoding experiences and data into symbolic representations conducive to formal reasoning, computation, and optimization of security policies. Recent works in West et al. (2022); Wu et al. (2023) provide promising approaches of using FMs to synthesize knowledge from datasets.
@@
Another challenging area of cyber deception is the configuration and utilization of honeypots in engaging attackers to acquire information and knowledge. Honeypot games have been used to formalize the deployment and configuration decisions by taking into account the adversarial interactions and the engagement goals. For instance, in Huang and Zhu (2019), a reinforcement learning algorithm is proposed to map out an attacker’s behavior. The engagement has to interact with the attacker, making sure that the attacker does not exist in the honeynetwork. By observing and analyzing attacker interactions with honeypots, defenders gain valuable insights into emerging threats and attacker behaviors. There is a tradeoff between learning the attacker’s behaviors and removing the attacker directly. The knowledge of attackers can be useful to further protect ourselves from zero-day attacks in the future. However, removing the attackers directly can lead to fewer risks.
@@ 4.2 A Neurosymbolic Learning Approach to Cyber Deception
FMs serve as pivotal tools across various functions including reinforcement learning, knowledge assimilation, formation of conjectures, and contextual representation. Moreover, they facilitate the coordination of multiple components in neurosymbolic learning. Figure 7 further illustrates the synergy between FMs and GMs for cyber deception. A sequence of tactical games $G_{1},G_{2},G_{3}$ unfold between an attacker and a defender. The defender leverages $F_{1}$ to conjecture the attacker’s behavior and utilizes $F_{2}$ to adapt security measures based on acquired observations. This conjecture and subsequent updates stem from the defender’s knowledge, which is symbolically represented using $F_{4}$.
@@
Figure 7: Interconnections between FMs and GMs for cyber deception: The tactical games are played between an attacker and a defender. A defender uses $F_{1}$ to conjecture the attacker’s behavior and uses $F_{2}$ to update the security knowledge. The conjecture and the update are based on the defender’s knowledge. The knowledge is symbolically represented and captured using $F_{4}$. The conjectured learning is driven by $F_{5}$, which coordinates the conjecture update and the policy update. The tactical game is sequentially played over time to achieve its operational-level objective. The changes in the environment and the operational-level objective can lead to the adaptation of tactical-level games. This change is represented by the contextual parameter $\theta$, which requires $F_{5}$ to process heterogeneous data sources to identify the context and $F_{6}$ to coordinate between the tactical level games and the contextual changes. Each context has an associated knowledge set $K$, which updates itself when the context changes. FM $F_{7}$ provides a set of rules for policy updates and conjecture formation under a context. FM $F_{4}$ is used to figure out the set of relevant knowledge which $F_{7}$ can choose from. $F_{4}$ can also create new rules for new contexts transferrable from other contexts or learned from datasets.
@@ 4.3 Emerging Challenges
Another practical challenge arises from the inference time current FMs consume to output the next token. In the presence of advanced adversarial attacks, and in particular, cyber threats, real-time inference is the prerequisite of a rapid defense response to counteract attackers’ swift moves, such as lateral movement within the system. Accelerating the FM’s inference with long contexts is imperative for predictive and prescriptive security analytics.